Given utility function U=X0.5Y0.5 where PX = 8 Birr, Birr, PY = 2 Birr and the income of the consumer is, M= 240 Birr.
a. Find the utility maximizing combinations of X and Y.
b. Calculate marginal rate of substitution of X for Y (MRSX,Y) at equilibrium and interpret your result.
1
Expert's answer
2020-10-05T13:09:30-0400

a. Finding the utility maximizing combinations of X and Y

  • Utility function

"U=X^{0.5}Y^{0.5}"

  • cost function

"240=8X+2Y"

  • Marginal Utility of X

"MU_{x}=\frac{dU}{dX}"

"MU_{x}=0.5X^{-0.5}Y^{0.5}"

  • Marginal Utility of Y

"MU_{y}=\frac{dU}{dY}"

"MU_{y}=0.5X^{0.5}Y^{-0.5}"

  • At equilibrium:

"MU_{x}=MU_{y}"

"0.5X^{-0.5}Y^{0.5}=0.5X^{0.5}Y^{-0.5}"

"0.5\frac{Y^{0.5}}{X^{0.5}}=0.5\frac{X^{0.5}}{Y^{0.5}}"

  • We may divide each side by 0.5 and then cross-multiply the equation

"{X^{0.5}}{X^{0.5}}={Y^{0.5}}{Y^{0.5}}"

"X=Y"

  • To solve the problem, let us substitute Y for X in the cost function

"240=8X+2Y"

"240=8X+2X"

"240=10X"

"X=24"

  • If "X=Y" then

"Y=24"


b. Calculating marginal rate of substitution of X for Y (MRSX,Y) at equilibrium and interpreting the result

"MRS_{xy}=\frac{MU_{x}}{MU_{y}}"

"MRS_{xy}=\frac{0.5X^{-0.5}Y^{0.5}}{0.5X^{0.5}Y^{-0.5}}" ="\frac{X^{-0.5}Y^{0.5}}{X^{0.5}Y^{-0.5}}" ="\frac{Y^{0.5}Y^{0.5}}{X^{0.5}X^{0.5}}"

"|MRS_{xy}|=\frac{Y}{X}"="\frac{24}{24}" =1

  • Interpretation:

The consumer is willing to give up one (1) unit of X to get an extra unit of Y and remain with a combination of goods that is equally satisfying